# AnomalyMatch: Discovering Rare Objects of Interest with Semi-supervised and Active Learning

### v1

@@ Abstract
Anomaly detection in large datasets is essential in fields such as astronomy and computer vision; however, supervised methods typically require extensive anomaly labelling, which is often impractical. We present AnomalyMatch, an anomaly detection framework combining the semi-supervised FixMatch algorithm using EfficientNet classifiers with active learning. By treating anomaly detection as a semi-supervised binary classification problem, we efficiently utilise limited labelled and abundant unlabelled images. We allow iterative model refinement in a user interface for expert verification of high-confidence anomalies and correction of false positives. Built for astronomical data, AnomalyMatch generalises readily to other domains facing similar data challenges. Evaluations on the GalaxyMNIST astronomical dataset and the miniImageNet natural-image benchmark under severe class imbalance (1% anomalies for miniImageNet) display strong performance: starting from five to ten labelled anomalies and after three active learning cycles, we achieve an average AUROC of $0.95$ (miniImageNet) and $0.86$ (GalaxyMNIST), with respective AUPRC of $0.77$ and $0.71$. After active learning cycles, anomalies are ranked with 71% (miniImageNet) to 93% precision in the 1% of the highest-ranked images. AnomalyMatch is tailored for large-scale applications, efficiently processing predictions for $\approx$ 100 million images within three days on a single GPU. Integrated into ESA’s Datalabs platform, AnomalyMatch facilitates targeted discovery of scientifically valuable anomalies in vast astronomical datasets. Our results underscore the exceptional utility and scalability of this approach for anomaly discovery, highlighting the value of specialised approaches for domains characterised by severe label scarcity.
Keywords Semi-supervised learning  $\cdot$ Active learning  $\cdot$ Anomaly detection  $\cdot$ High-performance computing
@@ 1 Introduction
Identifying anomalies – rare and unusual outliers – in large datasets is a central challenge across machine learning and applied sciences  [1, 2]. Anomalies often hold critical relevance, either as scientifically significant discoveries or as artefacts requiring removal.
There are various anomaly detection approaches, but they generally fall into three categories based on availability of labelled, i.e. known, anomalies: fully supervised, semi-supervised, and unsupervised methods. Fully supervised anomaly detection typically achieves the best performance but demands substantial labelled anomaly examples, which are often unavailable or prohibitively expensive to obtain  [3]. Conversely, purely unsupervised methods operate without labels but can struggle to discriminate scientifically meaningful anomalies from irrelevant outliers or noise and do not easily allow to make use of known information and anomalies  [4]. Semi-supervised learning (SSL) approaches, which use unlabelled data along with some limited labelled instances, offer a practical balance, leveraging minimal supervision to improve performance significantly over fully unsupervised approaches [3, 5].
Real-world anomaly detection workflows often benefit from directly incorporating human expertise via active learning and user-in-the-loop approaches [6]. These frameworks allow iterative model refinement through user interaction, crucially exploiting expert domain knowledge to identify the most valuable anomalies from a scientific perspective. The Astronomaly tool [7] is a prominent example in astronomy, effectively combining active learning with anomaly detection to iteratively refine its predictions based on user feedback. Additionally, citizen science projects such as Galaxy Zoo [8] have demonstrated the potential for non-expert volunteers to accelerate and scale the labelling process. Our work embraces this philosophy, emphasizing active learning to guide the anomaly detection pipeline toward scientifically interesting anomalies, rather than merely statistically unusual ones.
@@
In this work, we introduce AnomalyMatch, a semi-supervised anomaly detection tool that integrates FixMatch [5] with an EfficientNet-based neural architecture  [9], active learning, and a custom graphical user interface (GUI). This setup is then optimised for anomaly detection as a heavily imbalanced binary classification. Our framework enables domain experts or citizen scientists to iteratively label selected anomalies identified by the model during training, significantly improving detection performance through minimal additional labelling. Originally designed with astronomical datasets in mind, AnomalyMatch is highly generalizable to other domains with similar data characteristics. Moreover, our implementation prioritizes scalability: AnomalyMatch can efficiently process and analyse hundreds of millions of images on just one graphics card as demonstrated in the accompanying article  [10], making it particularly suitable for upcoming large-scale astronomical surveys and data repositories.
We validate AnomalyMatch using two datasets: miniImageNet (a natural-image benchmark) and GalaxyMNIST (an astronomy-specific galaxy morphology dataset). Our experiments quantify anomaly detection performance via area under the receiver operating characteristic (AUROC) and area under the precision recall curve (AUPRC) metrics under severe class imbalance as well as a anomaly detection efficiency metric. These benchmark results clearly demonstrate the efficacy of our semi-supervised and active-learning-based approach. Furthermore, AnomalyMatch complements existing methods effectively; while frameworks like PyOD [11] or Astronomaly [7] primarily rely on classical methods (e.g., Isolation Forest), AnomalyMatch combines semi-supervised and active learning with training a neural network . Thus, our method excels in targeted anomaly discovery, empowering users to define and detect anomalies of specific scientific interest rather than general statistical outliers.
@@
  • A novel semi-supervised anomaly detection pipeline (AnomalyMatch) combining FixMatch, EfficientNet, and active learning in a heavily optimised framework, effectively leveraging unlabelled data and greatly reducing labelling requirements.
@@
  • A GUI-based active learning loop enabling human experts (or citizen scientists) to iteratively guide anomaly selection and improve model performance during training.
@@
  • Comprehensive benchmarking of AnomalyMatch on both natural and astronomical image datasets, demonstrating excellent performance in terms of AUROC/AUPRC and anomaly detection efficiency with extremely limited labelled samples (five to ten labelled anomalies).
@@
  • A highly efficient and scalable implementation capable of analysing 100 million images within days  [10], seamlessly integrated into ESA’s Datalabs platform.
@@
  • A detailed discussion on implications for future large-scale anomaly detection applications, including potential expansions to new data modalities, advanced architectures (e.g., vision transformers), and explainable AI.
2 Related Work
2.1 Anomaly Detection Methods
Methods for anomaly detection, also known as outlier detection, are typically classified based on their dependence on labelled data into supervised, semi-supervised, and unsupervised learning paradigms [1].
Unsupervised methods identify anomalies based solely on deviations from the normal data distribution without labelled examples. Classical approaches include One-Class SVM [12], Isolation Forest [4], and Local Outlier Factor (LOF) [13], extensively available in libraries such as PyOD [11]. While effective on lower-dimensional or structured datasets, these methods often struggle with complex high-dimensional data like images, where feature extraction and similarity metrics are more challenging [14, 15].
Supervised methods, which require extensive labelled datasets containing anomalies, generally yield higher accuracy. However, fully labelled datasets are often impractical due to the difficulty and cost associated with obtaining labelled anomalies [3].
Semi-supervised methods, using a limited number of labelled examples and abundant unlabelled data, balance performance and practicality [16]. Deep learning-based semi-supervised anomaly detection approaches, such as Deep SAD [3], incorporate few labelled anomalies to guide learned representations. Similarly, Astronomaly [7] integrates active learning and semi-supervised learning principles into an astronomy-focused anomaly detection framework.
While classical approaches remain effective in simpler domains, neural networks have proven particularly promising for anomaly detection in complex, high-dimensional image data due to their inherent ability to learn rich, hierarchical representations from raw input [17].
2.2 Advances in Semi-Supervised Learning
General-purpose semi-supervised learning methods have rapidly advanced in recent years. Early SSL algorithms, including $\Pi$-Model, Pseudo-Label, VAT, Mean Teacher, and UDA, effectively leveraged unlabelled data through consistency regularization and pseudo-labelling [18, 19, 20]. Among these, FixMatch [5] has notably simplified the approach by combining consistency regularization and confidence-based pseudo-labelling. Specifically, FixMatch generates pseudo-labels from weakly augmented unlabelled data and strongly augmented versions and then trains the model to predict consistently on both, thus effectively making use of unlabelled data. This technique achieves state-of-the-art results on common benchmarks (e.g., CIFAR-10, SVHN) with minimal labelled examples.
Variants of FixMatch have been successfully applied to specific domains, such as remote sensing imagery [16], semantic segmentation [21, 22], and graph classification [23]. These adaptations highlight FixMatch’s flexibility in handling various data modalities and structured prediction tasks. The success of FixMatch and related approaches across diverse settings motivates their adaptation to anomaly detection tasks characterized by severe label scarcity and class imbalance. In this work, we leverage FixMatch by formulating anomaly detection as a two-class classification problem—“anomaly vs. normal”—thus fully exploiting the potential of SSL techniques in anomaly discovery.
Overall, our approach synthesizes state-of-the-art semi-supervised learning with domain-tailored active learning and scalable software implementation, bridging a methodological gap in anomaly detection efficacy, especially relevant to data-intensive scientific applications in domains such as astronomy.
@@ 3 Methods
Our anomaly detection system, AnomalyMatch, consists of a semi-supervised binary classifier (nominal vs. anomaly), using the FixMatch algorithm, combined with an active-learning loop and an intuitive guided user interface (GUI). We detail each component below, highlighting key design decisions.
@@ 3.1 FixMatch for Anomaly Detection
We base our training methodology on FixMatch [5], initially proposed for multi-class semi-supervised learning. We adapt it specifically to anomaly detection, treating the problem as a binary classification: “nominal" (regular data) vs. “anomaly" (rare, interesting data) with heavy class imbalance. We use a small labelled set $\mathcal{D}_{L}=\{(x_{i},y_{i})\}$ and a large unlabelled set $\mathcal{D}_{U}=\{u_{j}\}$, typically with $|\mathcal{D}_{U}|\gg|\mathcal{D}_{L}|$ and severe imbalance in favour of nominal examples. A practical difference to applying FixMatch in our scenario is that additional labels for nominal data are easier to obtain, whereas anomaly labels are often extremely scarce or difficult to expand due to their rarity. To handle this severe imbalance, we include oversampling using a weighted random sampler with occurrence-based sampling probabilities of anomalies / nominal samples during training.
Let $f_{\theta}(x)$ denote the neural network model with parameters $\theta$, outputting the probability $p(y=\text{anomaly}|x)$ for input $x$. In each training step, unlabelled images from $\mathcal{D}_{U}$ undergo weak augmentation (simple flips, crops) to generate pseudo-labels if the model’s predicted probabilities exceed a high-confidence threshold $\tau$ (e.g., 0.95). A corresponding strongly augmented image (applying substantial random transformations) is then fed into the network, enforcing consistency by training the model to match its predictions to the pseudo-labels.
The overall loss comprises supervised binary cross-entropy loss (on labelled data) and unsupervised consistency loss (on confident pseudo-labelled data):
$L=L_{\text{sup}}+\lambda\cdot L_{\text{unsup}}$
We set use $\lambda=1$ to emphasize unlabelled data contributions. Importantly, our experiments showed that training for fewer iterations (on the order of hundreds of iterations rather than tens of epochs until full convergence) is beneficial in the human-in-the-loop setup, as prolonged training sessions risk overfitting to the initial labels. This is explored in more detail in the Results Section.
@@ 3.2 EfficientNet Backbone
For computational efficiency and high accuracy, we chose EfficientNetLite0 [9], which achieves efficiency by systematically balancing depth, width, and resolution of the network through compound scaling, as our backbone convolutional neural network using input images resized to $224\times 224$. Smaller input sizes, such as $150\times 150$, were also successfully employed in our companion paper searching Hubble Space Telescope images  [10]. . The network is initialized using ImageNet pre-trained weights, fine-tuning all layers during FixMatch semi-supervised training. Weight decay ($\ell_{2}$ regularization) is included to prevent overfitting.
@@ 3.3 Active Learning Loop and User Interface
Integrating human expertise is central to our pipeline. After initial training, the active learning loop identifies unlabelled images with the highest predicted anomaly scores, presenting them to the user through a dedicated interface (Figure 2). Users verify and label anomalies, correcting false positives and confirming interesting candidates, thus efficiently refining the model.
Figure 2: AnomalyMatch active learning interface built with ipywidgets. Users can inspect and label model-flagged candidates, adjust visual settings (e.g., brightness, contrast, RGB channels), and monitor performance via receiver operating characteristic and precision-recall curves. Model and label states can be saved and reloaded across sessions.
@@
  • Image Processing Tools: Adjustments for brightness, contrast, unsharp masking, and channel-specific filters (RGB) to support precise labelling.
  • Performance Feedback: Immediate feedback through AUROC and AUPRC metrics if a held-out test set was defined initially.
  • Interactive Data Management: Users can load batches of unlabelled data, save session labels and models, and iteratively continue cycles of labelling and retraining.
This user-friendly interface leverages ipywidgets, facilitating interactive exploration within Jupyter notebooks. A prebuilt setup integrated within the European Space Agency’s (ESA) Datalabs platform further provides seamless access to large-scale astronomical datasets, such as the recent Euclid Q1 data release [24], enabling immediate practical use in scientific analyses.
3.4 Data Augmentations
Robust data augmentation is essential to effective semi-supervised learning. We started from augmentation strategies validated by MSMatch [16] and DistMSMatch [25] for the two-tiered augmentation approach:
  • Weak Augmentations: Minor transformations (horizontal flips, slight random crops) that preserve image content integrity, essential for accurate pseudo-label generation.
  • Strong Augmentations: Applying RandAugment [26], we perform aggressive transformations including rotations, colour distortions (brightness, contrast, solarization), sharpness variations, shearing, translations, and posterization. These diverse distortions encourage model robustness and invariance to image variability.

### v2

@@ Abstract
Anomaly detection in large datasets is essential in astronomy and computer vision. However, due to a scarcity of labelled data, it is often infeasible to apply supervised methods to anomaly detection. We present AnomalyMatch, an anomaly detection framework combining the semi-supervised FixMatch algorithm using EfficientNet classifiers with active learning. AnomalyMatch is tailored for large-scale applications and integrated into the ESA Datalabs science platform. In this method, we treat anomaly detection as a binary classification problem and efficiently utilise limited labelled and abundant unlabelled images for training. We enable active learning via a user interface for verification of high-confidence anomalies and correction of false positives. Evaluations on the GalaxyMNIST astronomical dataset and the miniImageNet natural-image benchmark under severe class imbalance display strong performance. Starting from five to ten labelled anomalies, we achieve an average AUROC of $0.96$ (miniImageNet) and $0.89$ (GalaxyMNIST), with respective AUPRC of $0.82$ and $0.77$. After three active learning cycles, anomalies are ranked with 76% (miniImageNet) to 94% (GalaxyMNIST) precision in the top 1% of the highest-ranking images by score. We compare to the established Astronomaly software on selected ’odd’ galaxies from the ’Galaxy Zoo - The Galaxy Challenge’ dataset, achieving comparable performance with an average AUROC of $0.83$. Our results underscore the exceptional utility and scalability of this approach for anomaly discovery, highlighting the value of specialised approaches for domains characterised by severe label scarcity.
keywords:
Semi-supervised learning – Active learning – Anomaly detection – High-performance computing
@@ 1 Introduction
Identifying anomalies – rare and unusual outliers – in large datasets is a central challenge in machine learning and the applied sciences  (Chandola et al., 2009; Pimentel et al., 2014). The term ‘anomaly’ has various definitions dependent upon the field they are being studied in. They can be telemetry outliers, instrumentation effects, objects exhibiting odd morphologies in images or outliers in well-measured distributions. Here we focus on searching for anomalies in image data. Some examples of astrophysical anomalies are gravitational lenses (Garvin et al., 2022; Euclid Collaboration et al., 2025), galaxy mergers (Pearson et al., 2022; O’Ryan et al., 2023), edge-on proto-planetary disks (Berghea et al., 2024), or ring galaxies (Abraham et al., 2025). Each of these objects provide unique insights into different aspects of astrophysics.
Anomalies can hold critical relevance as scientifically meaningful discoveries, whether aiding our understanding in galaxy evolution, cosmology or the evolution of objects within the Galaxy. By their nature, anomalies are rare and require large or deep surveys to be detected in large quantities. Next-generation observatories, including Euclid (Laureijs et al., 2011) and the Vera C. Rubin Observatory (Ivezić et al., 2019), are poised to release unprecedented volumes of astronomical data across an unmatched area of the sky. With such large data volumes on the horizon, we require efficient methodologies to search these colossal datasets for rare objects where labelled data are limited or, potentially, non-existent.
A key challenge in searching anomalies is that, given their rarity, there is typically a lack of labelled examples. Existing catalogues may consist of the order of tens of objects, for instance edge on protoplanetary disks (ALMA Partnership et al., 2015) or quadruply lensed quasars (Stern et al., 2021). We must also employ efficient techniques to identify them in datasets with the scale likely to contain them (i.e., sky spanning surveys such as Euclid). Due to the need for efficient processing of such datasets, machine learning methods are often employed for anomaly detection tasks.
Supervised machine learning methods are most commonly applied to classification tasks in astronomy. Many works have successfully used supervised machine learning to classify astrophysical sources based on their morphology, and expand labelled sets of objects (Walmsley et al., 2023). However, they are often ineffective without large labelled datasets to train them. A promising avenue has been to fine-tune pre-trained neural network models for anomaly detection tasks. Fine-tuning involves taking a model pre-trained on a large dataset and continuing to train it on a new task or different data - e.g. classifying a specific anomaly. This reduces the need for large labelled datasets and training from scratch, but a few hundred examples are often still needed - larger than available catalogues for many types of anomaly - and effectiveness may vary depending on transferability of features and concept drift.
Many works attempt to build the labelled datasets themselves to train supervised methods. Such approaches include tasking citizen scientists (e.g. Willett et al., 2013; Mantha et al., 2024) or have experts visually inspect a dataset to identify candidate anomalies (e.g. Acevedo Barroso et al., 2025). However, it has been found that special care must be taken when instructing individuals to search for different anomalies, or odd objects, as their definition can vary between different citizen scientists and experts (Mantha et al., 2024). Another issue is the size of observational datasets to search, as the scale of upcoming surveys makes visual inspection of a large portion of the data impossible.
An approach which can bypass the need for labelled data entirely is that of unsupervised machine learning methods. These methods do not require labelled data at all, and utilise the structure of the underlying data itself in order to identify different classes of objects (Schölkopf et al., 2001; Zhao et al., 2019). While effective on lower-dimensional or structured datasets, these methods often struggle with complex high-dimensional data like images, where feature extraction and similarity metrics are more challenging (Goldstein & Uchida, 2016; Zimek et al., 2012).
However, it is desirable to utilise the known anomalies, even if few, or annotated non-anomalous data when training a model. In many non-astronomical contexts, this has been achieved through semi-supervised learning techniques such as MSMatch (Gómez & Meoni, 2021), derived from the well-used FixMatch method (Sohn et al., 2020). These frameworks combine both supervised and unsupervised methods to learn about the data. MSMatch was able to achieve accuracy competitive with supervised methods on scene classification of satellite imagery with as few as five labelled examples per class while utilising thousands of unlabelled images.
Another approach, combining classical feature space exploration with an element of user-in-the-loop active learning, is Astronomaly (Lochner & Bassett, 2021). This algorithm successfully leverages the features of galaxies to break down datasets into different classes, which can then be directed by a user to identify objects of interest. Astronomaly was successfully benchmarked on the Galaxy Zoo: The Galaxy Challenge dataset, a soft label dataset derived from the Galaxy Zoo 2 (Willett et al., 2013) campaign. They found they were able to recover ‘Odd’ galaxies from a larger dataset with no pre-existing labelled data of the objects. This work showed that real-world anomaly detection workflows often benefit from directly incorporating human expertise via active learning and user-in-the-loop approaches (Settles, 2009). Such a framework allows iterative model refinement through user interaction, crucially exploiting expert domain knowledge to identify the most valuable anomalies from a scientific perspective. Thus, combining a semi-supervised method - leveraging labelled and unlabelled data - with a user-in-the-loop workflow appears to be key for furthering anomaly detection methods in astronomy.
@@
In this work, we introduce AnomalyMatch - a method that embraces this philosophy. AnomalyMatch is a semi-supervised anomaly detection tool that integrates FixMatch with an EfficientNet-based neural network architecture  (Tan & Le, 2019), active learning, and a custom graphical user interface (GUI). This setup is optimised for anomaly detection as a heavily imbalanced binary classification. Our framework enables domain experts or citizen scientists to iteratively label selected anomalies identified by the model during training, significantly improving detection performance through minimal additional labelling. Originally designed with astronomical datasets in mind, AnomalyMatch is highly generalisable to other domains with similar data characteristics. Moreover, our implementation prioritises scalability: AnomalyMatch can efficiently process and analyse hundreds of millions of images on just one graphics card, as demonstrated in the accompanying article (O’Ryan & Gómez, 2025), which makes it particularly suitable for upcoming large-scale astronomical surveys and data repositories. Thus, our method excels in targeted anomaly discovery, empowering users to define and detect anomalies of scientific interest rather than general statistical outliers.
@@
  • A novel semi-supervised anomaly detection pipeline (AnomalyMatch) combining FixMatch streamlined for heavy class imbalance, EfficientNet neural networks, and active learning in a heavily optimised framework, effectively leveraging unlabelled data and greatly reducing labelling requirements;
@@
  • A GUI-based active learning loop enabling experts to iteratively guide anomaly selection and improve model performance during training;
  • Seamless integration of established astronomy image processing tools, such as different normalisation methods from Astropy (Astropy Collaboration et al., 2013, 2018) and FITS file support.
@@
  • Comprehensive benchmarking of AnomalyMatch on both daily-life and astronomical image datasets, demonstrating excellent performance in terms of Area Under the Receiver Operating Characteristic (AUROC), Area Under the Precision-Recall Curve (AUPRC), and Anomaly Detection Efficiency with extremely limited labelled samples (five to ten labelled anomalies);
@@
  • A highly efficient and scalable implementation capable of analysing 100 million images within days  (O’Ryan & Gómez, 2025), seamlessly integrated into the European Space Agency’s (ESA) Datalabs platform [^1] https://datalabs.esa.int/ (Navarro et al., 2024);
@@
  • A detailed discussion on implications for future large-scale anomaly detection applications, including potential expansions to new data modalities, advanced architectures (e.g. vision transformers), and explainable AI.
Section 2 presents the components of the AnomalyMatch pipeline and provides the reasoning behind its key design choices. Section 3 describes the extensive testing and benchmarking of AnomalyMatch on multiple datasets and compares directly to the well used anomaly detection software Astronomaly on a dataset of anomalies from Galaxy Zoo 2. Section 4 discusses the significance of our anomaly detection framework and outlines its assumptions and limitations. Section 5 includes concluding remarks and discusses potential avenues for future work.
@@ 2 Methods
Our anomaly detection method, AnomalyMatch, consists of a semi-supervised binary classifier (normal vs. anomaly), using the FixMatch algorithm. We have combined this with an active-learning loop and an intuitive GUI. We detail each component here, highlighting key design decisions.
@@ 2.1 FixMatch for Anomaly Detection
We base our training methodology on FixMatch (Sohn et al., 2020), initially proposed for multi-class semi-supervised learning. FixMatch generates pseudo-labels from weakly and strongly augmented unlabelled data and then trains the model to predict consistently on both. Variants of FixMatch have been successfully applied to specific domains, such as remote sensing imagery (Gómez & Meoni, 2021), semantic segmentation (Martí i Rabadán et al., 2022; Yang et al., 2022), and graph classification (Jalali et al., 2024). These adaptations highlight FixMatch’s flexibility in handling various data types and prediction tasks. The success of FixMatch and related approaches across diverse settings motivates their adaptation to anomaly detection tasks characterised by severe label scarcity and class imbalance.
We adapt FixMatch to treat the problem of anomaly detection as a binary classification: "normal" (regular data) vs. "anomaly" (rare, interesting data) with heavy class imbalance. We use a labelled set $\mathcal{D}_{L}=\{(x_{i},y_{i})\}$ and an unlabelled set $\mathcal{D}_{U}=\{u_{j}\}$, typically with $|\mathcal{D}_{U}|\gg|\mathcal{D}_{L}|$ and severe imbalance in favour of normal examples. A practical difference to applying FixMatch in our scenario is that additional labels for normal data are easier to obtain, whereas the anomaly labelled dataset is small and difficult to expand due to their rarity. To handle this severe imbalance, we include oversampling using a weighted random sampler with occurrence-based sampling probabilities of anomalies / normal samples during training.
To train in an unsupervised manner, consistency regularisation is used. Let $f_{\theta}(x)$ denote the neural network model with parameters $\theta$, outputting the score $p(y=\text{anomaly}|x)$ for input $x$. In each training step, unlabelled images from $\mathcal{D}_{U}$ undergo weak augmentation (simple flips, crops) to generate pseudo-labels if the model’s predicted probabilities exceed a high-confidence threshold $\tau$ (e.g., 0.95). The same unlabelled images are strongly augmented (applying substantial random transformations) and then fed into the network, enforcing consistency by training the model to match its predictions to the pseudo-labels.
This only comprises half the loss function when training the model. The overall loss combines the unsupervised consistency regularisation with a supervised binary cross-entropy loss calculated using labelled data as
$L=L_{\text{sup}}+\lambda\cdot L_{\text{unsup}}.$ (1)
We set $\lambda=1$ in Equation 1 to emphasize unlabelled data contributions. Importantly, we find that training for fewer iterations (on the order of hundreds of iterations rather than tens of epochs) is beneficial in the human-in-the-loop setup, as prolonged training sessions risk overfitting to the initial labels. This is explored in more detail in the Section 3.
@@ 2.2 EfficientNet Backbone
For computational efficiency and high accuracy, we chose EfficientNetLite0 (Tan & Le, 2019). This architecture achieves improved efficiency by systematically balancing depth, width, and resolution of the network through compound scaling as our backbone convolutional neural network uses input images resized to $224\times 224$. Smaller input sizes, such as $150\times 150$, were also successfully employed in our companion paper searching Hubble Space Telescope images  (O’Ryan & Gómez, 2025). The network is initialised using ImageNet pre-trained weights, fine-tuning all layers during FixMatch semi-supervised training. Weight decay ($\ell_{2}$ regularisation) is included to prevent overfitting.
@@ 2.3 Active Learning Loop and User Interface
Leveraging human expertise is central to our pipeline. After initial training, the unlabelled data used is scored and ranked from most to least anomalous. A user is able to then confirm or correct the algorithm, both expanding the training set and refining the model. This also gives initial feedback to the user on the models performance. Figure 2 shows the dedicated interface through which the user conducts the active learning loop.
Figure 2: AnomalyMatch active learning interface built with ipywidgets. Users can inspect and label model-flagged candidates, adjust visual settings (e.g., brightness, contrast, RGB channels), and monitor performance via AUROC and AUPRC curves. Model and label states can be saved and reloaded across sessions.
@@
  • Image Display Tools. Adjustments for brightness, contrast, unsharp masking, and channel-specific filters (RGB) to support precise labelling;
  • Image Processing Tools. Adjustments for the normalisation and stretch applied when loading, training and evaluating images;
  • Performance Feedback. Immediate feedback through AUROC and AUPRC metrics if a test set was defined initially;
  • Interactive Data Management. Users can load batches of unlabelled data, save session labels and models, and iteratively continue cycles of labelling and retraining.
This user-friendly interface leverages ipywidgets, facilitating interactive exploration within Jupyter notebooks. A prebuilt setup integrated within ESA Datalabs further provides seamless access to large-scale astronomical datasets, such as the Euclid Quick Data Release (Q1; Aussel et al., 2025). This enables immediate practical use in scientific analyses.
2.4 Data Processing and Augmentations
Robust data augmentation is essential to effective semi-supervised learning. We start from augmentation strategies validated by MSMatch (Gómez & Meoni, 2021) and DistMSMatch (Östman et al., 2023) for the two-tiered augmentation approach:
  • Weak Augmentations. Minor transformations (horizontal flips, slight random crops) that preserve image content integrity, essential for accurate pseudo-label generation;
  • Strong Augmentations. Applying RandAugment (Cubuk et al., 2020), we perform aggressive transformations including rotations, colour distortions (brightness, contrast, solarisation), sharpness variations, shearing, translations, and posterisation. These diverse distortions encourage model robustness and invariance to image variability.